WHAT IS FOURRIER SERIES
PLEASE EXPLAIN FROM THE BASICS 

Grade:12

1 Answers

sudhir pal
10 years ago
expression of this type are fourier polynomial
\begin{displaymath}F_n(x) = a_0 + \sum_{k= 1}^{k=n} \Big(a_k\cos(kx) + b_k\sin(kx)\Big).\end{displaymath}
fourier polynomial has period (2pi/k)
fourier series is summation of infinite fourier polynomial
\begin{displaymath}A_0 + \sum_{n = 1}^{\infty} (A_n\cos(nx) + B_n\sin(nx)).\end{displaymath}
where A0 is constant and An & Bn are cofficients
this question is not important for those who are preparing for IIT-JEE and other engineering entrance exams in INDIA